# Sample Compression for Continual Learning

### v1

@@ 2.2 Sample Compression Theory
2.3 Pick-to-Learn
To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM (Boser et al., 1992) and the perceptron (Rosenblatt, 1958; Moran et al., 2020), have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) (Marchand & Shawe-Taylor, 2002; Marchand et al., 2003; Marchand & Sokolova, 2005) and decision trees (Shah, 2007), necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan et al. (2024) as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm modifies the training loop of the predictor, by iteratively choosing datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of Pick-To-Learn in Algorithm 1.
Algorithm 1 Original Pick-To-Learn (P2L)
@@ 2.3 Pick-to-Learn
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
@@
9:  return $\theta_{k},C_{k}$.
10:  {Learned parameters and compression set}
@@ 3 Sample Compression for Continual Learning
3.1 A New Training Scheme
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer, thus mitigating the forgetting by adding a well-chosen datapoint.
@@ 3.1 A New Training Scheme
0:  $\theta_{\rm init}$Initialization parameters of a given model $f_{\theta}$}
@@
11:     $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$
@@
13:  return $\theta_{k},C_{k}$.
14:  {Learned parameters and compression set}
@@
0:  $\theta_{0}$Initialization parameters of a given model}
0:  $\gamma$P2L’s stopping criteria}
@@
6:     $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\quad\forall i=1,\ldots,t-1$
@@
10:  return $\theta_{T}$. {Learned parameters}
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm CoP2L calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum size of the buffer. We present CoP2L in Algorithm 3.
We modified Pick-To-Learn in two significant ways. First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer.
The second modification leading to mP2L is the choice of the returned predictor. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior. Instead, to perform early stopping, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
@@
From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) the bound function $\Psi(S,\theta,C)$ cannot be computed on the whole dataset, when only the compression set is given as input; (2) the sampling procedure sample$(S,m)$ cannot be reproduced in the reconstruction step. In the following section, we define the compression and reconstruction functions of CoP2L, which will give rise to sample compression bounds as presented in Theorem 3.1.
3.2 Compression and Reconstruction Scheme
To obtain generalization bounds for CoP2L, we need to prove that it can be compressed. To do so, we provide the compression and reconstruction functions of CoP2L. The compression function is presented in Algorithm 4 (provided in supplementary material) and the reconstruction function is presented in Algorithm 5.
Compression function.

### v3

@@ 2 Background
Pick-to-Learn.  To obtain sample compression guarantees for a class of predictors, it is necessary to prove that the learned predictor only depends on a small subset of the data and (optionally) a message. Some predictors, such as the SVM boser_svm and the perceptron rosenblatt1958perceptron ; moran2020perceptron , have straightforward compression scheme, as they only depend on a subset of the data after training. Some algorithms, such as the set covering machine (SCM) marchand2002set ; marchand2003set ; marchand2005learning and decision trees shah_sample_2007 , necessitate hand-crafted compression schemes involving a message. Up until recently, there was no sample compression scheme for deep neural networks.
Algorithm 1 Original Pick-To-Learn (P2L) paccagnan_pick_learn_2023
@@
0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
@@
9:  return $\theta_{k}$Final parameters}
The meta-algorithm Pick-To-Learn (P2L) was proposed by Paccagnan and Campi paccagnan_pick_learn_2023 as a compression scheme for any class of predictors, with a specific focus to deep neural networks. The meta-algorithm acts as an aditional training loop, by iteratively selecting datapoints over which the model is updated. Starting with initial parameters $\theta_{\rm init}$, P2L evaluates the predictor on the whole training dataset, adds the datapoints with the highest losses to the compression set and then updates the predictor using the compression set. The meta-algorithm stops once the losses of training examples not in the compression set are lower than a given stopping criteria ($-\ln(0.5)$ for the cross-entropy loss, which is equivalent to achieving zero errors on the complement set $S^{\operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}}$). We provide the pseudo-code of the original P2L in Algorithm 1. Recent works introduce improvements to the original scheme, notably early stopping strategies marks2025pick2learn-gp ; bazinet2024 and extension to the Bayesian setting marks2025pick2learn-gp .
@@ 3 Sample Compression for Continual Learning
A striking realization coming from the P2L algorithm is that only a small fraction of the training set – the compression set – needs to be provided to the learner in order to achieve good generalization. The winning strategy is to select this compression set to ensure a low risk on the training samples not being retained in the compression set (i.e., the complement of the compression set). This observation motivates our strategy for managing the replay buffer of our new algorithm Continual Pick-To-Learn (CoP2L): we subsample the complement set to create the replay buffer, in order to maintain a low risk on these examples while learning subsequent tasks. When needed, CoP2L can choose to add a datapoint from the buffer to the compression set, thus mitigating the forgetting by focusing on a well-chosen datapoint.
@@
0:  $\theta_{\rm init}$Initial parameters of the model $f_{\theta}$}
@@
11:    $k\leftarrow\operatorname*{\mathrm{argmin}}_{0\leq k^{\prime}\leq k}\Psi(S,\theta_{k^{\prime}},C_{k^{\prime}}\cap S)$Eq. (6)}
@@
13:  return $\theta_{k},C_{k}$. {Final parameters and compression set}
@@
0:  $\gamma$mP2L’s stopping criteria}
@@
6:    $B_{i}\leftarrow\mbox{sample}(B_{i},\lfloor\frac{m}{t}\rfloor)\ \forall i{=}1,\ldots,t{-}1$
@@
10:  return $\theta_{T}$. {Final parameters}
Each new task is learned by a modified version of the P2L meta-algorithm (Algorithm 2, entitled mP2L), as explained further down. That is, for each new task, the proposed continual learning algorithm (Algorithm 3, entitled CoP2L) calls mP2L and updates the buffer using the datapoints that were not chosen in the compression set. Similarly to the replay buffer method, at the end of each task $t$, the buffer contains $\lfloor\tfrac{m}{t}\rfloor$ datapoints of each previous task, with $m$ the maximum buffer size.
We modified Pick-To-Learn in two significant ways. (i) First, we introduced weights to the loss functions to tackle the imbalance problem between the current task and the previous tasks. When working with a replay buffer, the class imbalance problem is taken care of by training on a number of datapoints from previous tasks. As our model is only trained on a very limited subset of the data, we need to find another way to mitigate the effect of class imbalance. Before starting the training on a new task, we set the weight of the datapoints from the previous task to $\omega$ and the weight of the current task to $1$. For the stopping criteria of mP2L to be satisfied, the worst loss on the current task must be less than $\gamma$ and the worst loss on the previous task must be less than $\tfrac{\gamma}{\omega}$. This ensures that the model has achieved perfect accuracy on the current task and all the datapoints from the previous tasks that are found in the buffer. (ii) The second modification is the choice of the returned predictor by mP2L. The original P2L algorithm trains the model until it first achieves zero errors on the complement set. In presence of a noisy dataset, this might lead to an overfitting behavior, P2L adding most training samples into its compression set, which in turn gives a large generalization bound value. Instead, we perform early stopping based on the bound value (as proposed by bazinet2024 ). That is, mP2L relies on the trade-off encoded in Theorem 2.1 between the accuracy on the complement set and the complexity of the model. More precisely, it returns the model’s parameters $\theta$ that minimizes
@@
Compression and reconstruction scheme.  From a theoretical standpoint, in contrast to the original P2L, both mP2L and CoP2L cannot be reconstructed straightforwardly without a message for two reasons: (1) when only the compression set is given as input, the bound function $\Psi(S,\theta,C)$ used as stopping criteria in mP2L cannot be computed on the whole dataset; (2) the sampling procedure sample$(S,m)$ in CoP2L cannot be reproduced in the reconstruction step. In Appendix A, we provide the compression and reconstruction functions of CoP2L, which give rise to sample compression bounds as presented in Theorem 3.1. Noteworthy, after learning on $T$ tasks, the compression function provides two sample compression sets $S^{\operatorname{\mathbf{i}}}$ and $S^{\operatorname{\mathbf{j}}}$, along with a message pair $(\mu_{1},\mu_{2})$ chosen among the set of all possible messages, denoted as $\mathscr{M}(\operatorname{\mathbf{j}},T)=\{2,\ldots,T\}^{|\operatorname{\mathbf{j}}|}\times\left[\boldsymbol{\times}_{t=1}^{T}\{1,\ldots,n_{t}+|{\mathcal{B}}|\}\right].$ As explained further in appendix, the use of two compression sets, along with the message $\mu_{1}=[\mu_{1}^{i}]_{i=1}^{T}$, serves to the proper treatement of the buffer $B^{\star}$, while the message $\mu_{2}=[\mu_{2}^{i}]_{i=1}^{T}$ contains the number of iterations $K^{\star}$ to perform for each call to mP2L.
Generalization bound.  Given the compression and reconstruction function detailed in Apprendix A, Theorem 2.1 can be used to obtain generalization bounds on the last learned task, which is done by mP2L to compute the bound $\Psi$. However, this bound holds for only one distribution of data. Therefore, it cannot be applied to bound the risks of the learned predictor on previously tasks seen by CoP2L. The next Theorem 3.1 holds for any previously learned tasks. For a dataset $S_{t}\sim\operatorname{\mathcal{D}}_{t}$, we denote the loss on the complement set of the joint set $(S_{t}^{\operatorname{\mathbf{i}}}\cup S_{t}^{\operatorname{\mathbf{j}}})^{\rm\bf c}$ as ^$\widehat{{\mathcal{L}}}_{S_{t}}^{\ \operatorname{\operatorname{\mathbf{i}}^{{\rm\bf c}}}\cap\operatorname{\operatorname{\mathbf{j}}^{{\rm\bf c}}}}\!(\theta)$. Moreover, we denote the reconstruction function of CoP2L as $\operatorname{\mathscr{R}}_{1:T}(S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu)=\operatorname{\mathscr{R}}\big{(}S_{t}^{\operatorname{\mathbf{i}}},S_{t}^{\operatorname{\mathbf{j}}},\mu\mid S_{1},\ldots,S_{t-1},S_{t+1},\ldots,S_{T}\big{)}.$ This formulation is important for the following theorem, as the reconstruction function for task $t$ is conditioned on all datasets $S_{1}$ to $S_{T}$, with the exception of $S_{t}$.
Theorem 3.1.
$\forall t\in\{1,\ldots,T\},{\mathbf{i}}\in\operatorname{\mathscr{P}}(n_{t}),\operatorname{\mathbf{j}}\in\operatorname{\mathscr{P}}(n_{t}),\mu=(\mu_{1},[\mu_{2}^{i}]_{i=1}^{T})\in\mathscr{M}(\operatorname{\mathbf{j}},T):$
$\leq\tfrac{1}{n_{t}-|{\mathbf{i}}|-|\mathbf{j}|}\big{[}\log{\textstyle{n_{t}\choose|{\mathbf{i}}|}}+\log{\textstyle{n_{t}-\operatorname{|\mathbf{i}|}\choose|\mathbf{j}|}}+\textstyle\sum_{i=1}^{T}\log\tfrac{1}{\zeta(\mu_{2}^{i})}+|\mathbf{j}|\log(T-1)+\log\tfrac{T}{\zeta(|{\mathbf{i}}|)\zeta(|\operatorname{\mathbf{j}}|)\delta}\big{]}\,.$
The proof is given in Appendix C. It first relies on a tighter version of Theorem 2.1, where the $2\sqrt{n-\operatorname{|\mathbf{i}|}}$ term is removed and a second compression set is added. Then, this new result is adapted to the continual learning setting. For all tasks $1$ through $T$, this new theorem encodes a tradeoff similar to Theorem 2.1 between the predictor accuracy and its complexity. Here, the predictor complexity is evaluated using the size of the compression sets $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ and the probability of choosing a message. The probability of the messages $\mu_{1}$ and $\mu_{2}$ are functions of the size of $\operatorname{\mathbf{j}}$ and the number of tasks. Indeed, the probability of $\mu_{1}$ decays when $|\mathbf{j}|$ and $T$ grow larger. Moreover, the probability of $\mu_{2}$ depends on the size of the compression set outputted by mP2L at each task $T$. Thus, the larger the compression sets, the smaller the probability of both messages. In conclusion, the bound encodes the ability of a model to minimizes the loss on the dataset whilst still keeping $\operatorname{\mathbf{i}}$ and $\operatorname{\mathbf{j}}$ the small.
4 Experiments
Datasets and continual learning settings.  We carry-out experiments on MNIST lecun2010mnist , Fashion-MNIST fashionmnist , EMNIST (characters) cohen2017emnistextensionmnisthandwritten , PermutedMNIST goodfellow2013empirical , Rotated MNIST, CIFAR10 cifar10cifar100 , CIFAR100 cifar10cifar100 and TinyImageNet tiny-imagenet . We report results that include Class-Incremental (CI), Task-Incremental (TI), and Domain-Incremental (DI) continual learning settings. The dataset stats regarding how many classes/data items are available per task is given in Appendix H.
Baselines.  To assess the effectiveness of our proposed method CoP2L, we benchmark against several strong and widely adopted baselines across Class-Incremental, Task-Incremental, and Domain-Incremental continual learning settings. For each continual learning set-up we include baselines that yield the best performance in each category according to the Avalanche benchmark lomonaco2021avalanche .
For the Class-Incremental setting, we use the baselines Replay rolnick2019experience , iCaRL rebuffi2017icarl , and GDumb prabhu2020gdumb . Replay in particular is considered a strong baseline DellaLibera2023CLMASRAC due to its simplicity and effectiveness. In the Task-Incremental setting, we compare against Replay rolnick2019experience , LaMAML gupta2020look and Learning without Forgetting (LwF) li2017learning . These methods are well suited for TI scenarios where task identity is known at test time. In the Domain-Incremental setting, we evaluate against Replay rolnick2019experience , Less Forgetful Learning (LFL) jung2018less , Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , and Synaptic Intelligence (SI) zenke2017continual .
Implementation.  All our experiments are conducted using the Avalanche continual learning framework avalanchelib , which provides a unified interface for datasets, benchmarks, and baseline implementations. We also use the official ContinualAI benchmark repository [^1] https://github.com/ContinualAI/continual-learning-baselines for consistent implementation and fair comparisons. The code for the experimental results is available in the supplemental material.
Model architecture and training details.  We use the SimpleMLP architecture and customized SimpleCNN architectures (SimpleMLP and SimpleCNN from the Avalanche toolkit) for the MNIST, FMNIST, and EMNIST experiments. All models are trained using the SGD optimizer. The details of the models architectures and the hyperparameters used can be found in Appendix G. For experiments on more complex datasets such as CIFAR10, CIFAR100, and TinyImageNet, we use the pretrained feature extractor Dino caron2021emerging using ViT dosovitskiy2020image and ResNet50 he2016deep backbones, followed by a trainable MLP head. This choice is motivated by the insight that CoP2L can further benefit from stronger feature representations, especially in settings with limited data per class or high dataset complexity. By leveraging Dino’s obtained embeddings, we enhance the quality of the input representations, which we empirically observe to yield good performance. Results with standard deviations across seeds and total training times are reported in Appendix F. We used a single NVIDIA 4090 GPU for our experiments.
Accuracy and Forgetting.  We denote the accuracy obtained on a task $t$, after training the model for $T$ tasks with $\text{Accuracy}_{T,t}$. The accuracy we report is defined as,
. $\textit{Average Accuracy}_{T}=\frac{1}{T}\sum_{t=1}^{T}\text{Accuracy}_{T,t}$, where $t$ is for the task index. .
We also report the average forgetting at task $T$ which is defined as
$\textit{Average Forgetting}_{T}=\frac{1}{(T-1)}\sum_{t=1}^{T-1}[\text{Accuracy}_{t,t}-\text{Accuracy}_{T,t}],$ which is calculated when $T>1$.
4.1 Main Results
In Table 1, we present the average accuracy and average forgetting values defined in the section above, obtained on MNIST, FMNIST and EMNIST dataset, both with an MLP and a CNN architecture, where the networks are trained from scratch. In Appendix F, we also provide the standard deviations as well. We observe that with the MLP architecture CoP2L is able to outperform the baselines. Furthermore, we also observe that CoP2L is able to obtain competitive results when employed with a CNN architecture as well. In Appendix E, we also provide an analysis of the performance of CoP2L for varying buffer sizes.
In Table 2, we extended our results to cover CIFAR10, CIFAR100, and TinyImageNet datasets, for both Class-Incremental and Task-Incremental settings. We also provide results for Domain-Incremental setting with Permuted-MNIST goodfellow2013empirical and Rotated-MNIST datasets.
Table 1: Summary of average accuracies and forgetting values obtained on MNIST, FMNIST, EMNIST (5 tasks, Class-Incremental), and on Permuted-MNIST and Rotated-MNIST (10 tasks, Domain-Incremental).
Method / Dataset MNIST FMNIST EMNIST
Acc / For Acc / For Acc / For
CoP2L (MLP) 95.1 / 3.4 84.2 / 12.1 78.0 / 16.7
Replay (MLP) 94.5 / 6.1 84.0 / 15.9 77.2 / 22.1
GDumb (MLP) 91.7 / 3.8 81.8 / 10.4 62.6 / 16.0
CoP2L (CNN) 96.6 / 2.2 87.4 / 11.1 87.9 / 8.4
Replay (CNN) 96.9 / 3.3 87.6 / 13.3 87.7 / 12.3
GDumb (CNN) 94.4 / 3.0 85.0 / 8.8 85.1 / 7.5
Method / Dataset PermMNIST RotMNIST
Acc / For Acc / For
CoP2L (MLP) 92.3 / 6.0 88.3 / 9.8
Replay (MLP) 91.5 / 6.4 88.2 / 10.0
LFL (MLP) 83.5 / 9.3 51.6 / 45.7
EWC (MLP) 83.7 / 5.2 56.5 / 35.4
SI (MLP) 86.7 / 10.7 57.7 / 43.9
Table 2: Summary of average accuracies and forgetting values obtained on various continual learning tasks on CIFAR10, CIFAR100 and TinyImageNet datasets. CI denotes Class-Incremental, TI denotes Task-Incremental continual learning settings.
CIFAR10 (5 tasks), CI CIFAR100 (10 tasks), CI CIFAR100 (20 tasks), CI TinyImNet (40 tasks), CI
Method Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$) Acc ($\uparrow$) Forg ($\downarrow$)
CoP2L-Dino (ViT) 94.45 3.36 73.83 16.76 72.50 16.96 51.13 36.81
Replay-Dino (ViT) 94.01 6.01 68.51 30.44 69.23 29.60 47.66 48.15
iCaRL-Dino (ViT) 93.24 1.55 61.50 14.53 64.11 15.25 20.78 52.95
GDumb-Dino (ViT) 94.15 3.39 72.73 11.70 72.05 11.38 54.53 14.79
CoP2L-Dino (RN50) 81.13 5.55 49.18 21.99 49.98 24.32 32.74 36.40
Replay-Dino (RN50) 81.83 17.09 49.40 42.07 49.91 44.41 28.43 63.26
iCaRL-Dino (RN50) 74.56 1.54 45.97 8.71 46.87 6.32 33.13 10.95
GDumb-Dino (RN50) 80.70 8.95 43.83 20.31 43.78 19.03 24.04 21.82
CIFAR10 (5 tasks), TI CIFAR100 (10 tasks), TI CIFAR100 (20 tasks), TI TinyImNet (40 tasks), TI
CoP2L-Dino (ViT) 99.01 -0.03 95.15 0.92 97.10 0.71 94.28 1.22
Replay-Dino (ViT) 99.24 -0.03 96.59 0.23 98.06 0.41 95.82 0.36
LaMAML-Dino (ViT) 99.18 0.06 95.77 0.40 97.70 0.33 94.90 1.15
LwF-Dino (ViT) 97.53 2.08 95.26 1.06 95.73 2.15 85.79 9.09
CoP2L-Dino (RN50) 96.17 0.79 86.58 2.26 90.88 2.37 88.75 2.13
Replay-Dino (RN50) 96.96 0.42 87.59 0.64 93.13 0.16 90.53 0.74
LaMAML-Dino (RN50) 96.71 0.66 83.13 6.81 90.57 3.35 87.25 4.55
LwF-Dino (RN50) 95.90 0.15 79.29 0.67 87.94 0.09 85.15 0.90
Figure 1: Example demonstration of emprical upper bounds on test errors, on Class-Incremental (top-left CIFAR10, top-right CIFAR100) and Task-Incremental (bottom-left CIFAR10, bottom-right CIFAR100) datasets.
Empirical study of the CoP2L generalization bound.  In Figure 1, we showcase the empirical upper bounds on test errors. The details on the computation of the bound can be found in the Appendix  A. We observe that the bound values are non-trivial and generally follow similar trends as the actual error we report on the test set, while the estimate for the bound is calculated exclusively on the training set. In Figure 1, we provide plots for the bounds for the models trained with a buffer size of 2000, for Class-Incremental and Task-Incremental settings. CoP2L achieves particularly tight generalization bounds in the Task-Incremental setting, where forgetting is inherently reduced due to access to task identities and the use of a separate classifier head for each task. These conditions allow the model to better compress task-specific knowledge, leading to both stronger theoretical guarantees and improved empirical performance. We also observe that with the ViT backbone, the bounds are significantly tighter than those obtained with the ResNet50 backbone. This is likely because the ViT backbone provides more structured and information-rich representations, which are easier to compress effectively, thereby enhancing the performance of CoP2L.
We also give the results for all buffer sizes for the three MNIST datasets, and also bound estimates for the CIFAR10, CIFAR100, and ImageNet datasets (TinyImageNet tiny-imagenet , ImageNette imagenette and ImageWoof Howard_Imagewoof_2019 ) as well as bounds for the Domain-Incremental setting in Appendix D.
4.2 Analysis of CoP2L
In this section we provide an analysis of CoP2L with an ablation study on loss terms, a sensitivity analysis with respect to task ordering, and buffer size. Due to space limitations the figures are provided in the appendix, however the main results are mentioned here in the text.
Ablation on CoP2L loss terms  We have conducted an ablation study to understand the effect of additional weighting that counteracts class imbalance that arises in the continual learning setup when new tasks are introduced (Standard replay counteracts this by extending the batch size to have equal class representation). In the modified P2L (Algorithm 2), we incorporate an additional weighting term over classes to adapt P2L for class-balanced continual learning. In order to assess whether this idea is indeed helpful, we have run an experiment on the MNIST dataset with and without the additional weighting idea. In Appendix E, in Figure 10, we show this experiment. For both accuracy (left panel) and forgetting (right panel), the inclusion of the additional loss weighting term significantly improves the results.
[…]
Analysis on buffer size.  In Appendix E, in Figure 12, we compare the accuracy and forgetting obtained with different replay buffer sizes of (1000, 2000, 3000, 4000, 5000) samples for CoP2L and replay. This represents roughly 8% to 40% of the per-task dataset size, for MNIST and Fashion-MNIST, and 10% to 52% for EMNIST. For each replay buffer size, we report the variation of accuracy obtained at the end of training averaged over different seeds, and test accuracy for different tasks.
For the MNIST dataset, on average CoP2L significantly outperforms replay (as evidenced by the variation over seeds). For the Fashion-MNIST and EMNIST datasets, when using small buffer sizes, the average accuracy obtained with our CoP2L is significantly better. For larger buffer sizes, in general CoP2L performs comparably or slightly better. We also note for all three datasets, CoP2L significantly outperforms replay in terms of forgetting. We attribute this due to the fact that CoP2L is able to intelligently select examples from the replay buffer. Namely, CoP2L is able to focus on examples that are most representative of the earlier tasks, which we observe to help significantly in mitigating the deterioration over earlier tasks as evidenced by the results presented in the right column in Figure 12.
5 Limitations and Broader Impact Statement
We observe with our experiments that the performance of CoP2L is dependent to the ability of the feature extractor (backbone). In case a weak backbone is employed, CoP2L could necessitate a larger compression set (e.g. more data points per task). We have also seen that in those cases CoP2L might take longer to converge. We noted in our experiments that using deep pretrained networks such as ViT results in good performance, by increasing the compression ability. More broadly, we do not believe there is any particular negative social impact to note for our proposed methodology. On the contrary, we believe that our algorithm could be instrumental for improving trustworthiness of continual learning / lifelong learning in practice.
6 Conclusions
In this paper, we have proposed CoP2L, an algorithm rooted in the sample compression theory for self-certified continual learning. To the best of our knowledge, this is the first attempt in the literature to employ sample compression theory within the continual learning context. We provided sample compression bounds for CoP2L and verified empirically that they are non-vacuous and informative. We furthermore showed that on several challenging datasets, that our approach yields comparable or better results compared several strong baselines, while also providing learning guarantees, for trustworthy continual learning. As the combination of Pick-to-Learn with experience replay leads to a successful learning scheme, we foresee that combining Pick-to-Learn with other continual learning approaches could also lead to new compelling ways of obtaining self-certified predictors.
Acknowledgements
Mathieu Bazinet is supported by a FRQNT B2X scholarship (343192, doi: https://doi.org/10.69777/343192). Pascal Germain, and Cem Subakan are supported by Discovery Grants RGPIN-2020-07223 and RGPIN 2023-05759, respectively. This research was also enabled in part by support provided by Calcul Québec (https://www.calculquebec.ca/) and the Digital Research Alliance of Canada (https://alliancecan.ca/en).
Appendix A Compression and Reconstruction Functions